OpenStudy (anonymous):

What is the new volume of a rectangular prism that is originally 18 cubic meters and is then resized using a scale factor of 2 on each side length? a]9m ( to the 3rd power) b]36m ( to the 3rd power) c]72m ( to the 3rd power) d]144m ( to the 3rd power)

OpenStudy (anonymous):

18 resized using a scale factor of 2 18*2= 36 V=18^3 is b

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

18 resized using a scale factor of 2 18*2= 36 V=36^3 is b sorry 36 not 18 accident I type wrong

OpenStudy (anonymous):

36